# The Identity Problem in the special affine group of $\mathbb{Z}^{2}$

### v5

@@ 5.6. $H$ is generated by a positive scale $A$
    Hence, in both cases, $\boldsymbol{c}_{1}+C_{1}\boldsymbol{c}_{M}\in\operatorname{Cell}(\boldsymbol{d}_{ij})\subseteq\mathcal{S}$. We then show that $\boldsymbol{c}_{1}+C_{1}\boldsymbol{c}_{M}+C_{1}\boldsymbol{x}^{\prime}$ is a $\mathbb{R}_{>0}$-combination of elements in $\mathcal{S}$. Since $\boldsymbol{x}^{\prime}$ is either zero or a $\mathbb{R}_{>0}$-combination of elements in $\mathcal{S}$, write $\boldsymbol{x}^{\prime}=\sum_{i=1}^{m}r_{i}\boldsymbol{s}_{i}$, where $m\geq 0,r_{i}>0,\boldsymbol{s}_{i}\in\mathcal{S}$. Then $C_{1}\boldsymbol{x}^{\prime}=\sum_{i=1}^{m}r_{i}C_{1}\boldsymbol{s}_{i}$ is still a $\mathbb{R}_{>0}$-combination of elements in $\mathcal{S}$ by Equation (8). Hence, $\boldsymbol{c}_{1}+C_{1}\boldsymbol{c}_{M}+C_{1}\boldsymbol{x}^{\prime}$ is a $\mathbb{R}_{>0}$-combination of elements in $\mathcal{S}$.
  3. 3.
    If $t_{1}t_{M}=0$. Suppose $t_{1}=0$, the case where $t_{M}=0$ can be done analogously.
@@
    where $\boldsymbol{y}$ and $\boldsymbol{y}^{\prime}$ can be written as a $\mathbb{R}_{>0}$-linear combination of elements in $\mathcal{S}$. Therefore $\boldsymbol{x}=\boldsymbol{y}+\boldsymbol{y}^{\prime}$ also satisfies this claim.
@@
The semigroup $\langle\mathcal{G}\rangle$ is a group if and only if there exists a finite set of vectors $\{\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{m}\}\subseteq\mathcal{S}$ that satisfies the four conditions (i)-(iv) in Lemma 23.
@@
Therefore, $\boldsymbol{s}_{t}$ can always be written as a strictly positive linear combination of vectors $\boldsymbol{v}$ with $(I,\boldsymbol{v})\in\langle\mathcal{G}^{\prime}\rangle$. Hence, by condition (i), there exist strictly positive reals $r_{1},\ldots,r_{M}$ such that $r_{1}\boldsymbol{v}^{\prime}_{1}+\cdots+r_{M}\boldsymbol{v}^{\prime}_{M}=\boldsymbol{0}$, where $(I,\boldsymbol{v}^{\prime}_{t})\in\langle\mathcal{G}^{\prime}\rangle,t=1,\ldots,M$. Furthermore, conditions (ii), (iii) and (iv) show that every letter $(A^{\prime}_{i},\boldsymbol{a}^{\prime}_{i})$ in $\mathcal{G}^{\prime}$ appears at least once in a word representing $(I,\boldsymbol{v}^{\prime}_{t})$ for some $t\in\{1,\ldots,M\}$.
@@
Next, suppose $\langle\mathcal{G}\rangle$ is a group, we show that there exists an $\mathbb{R}_{>0}$-linear combination of elements in $\mathcal{S}$ equal to $\boldsymbol{0}$, that satisfies the conditions (ii), (iii) and (iv).
@@
Then by Lemma 23, the vector $\boldsymbol{0}$ can be written as a $\mathbb{R}_{>0}$-linear combination of a finite number of elements in $\mathcal{S}$, satisfying the conditions (ii), (iii) and (iv).
@@
If the generator $A$ of the group $\langle A_{1},\ldots,A_{K}\rangle\cong\mathbb{Z}$ is a positive scale, it is the decidable in PTIME whether $\langle\mathcal{G}\rangle$ is a group.
@@
Since $\langle A_{1},\ldots,A_{K}\rangle\cong\mathbb{Z}$, we have $J_{+}\neq\emptyset,J_{-}\neq\emptyset$. Given $\mathcal{S}$ as a set of cells, it is decidable whether there exist a finite non-empty set of vectors $\{\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{m}\}\subseteq\mathcal{S}$ satisfying conditions (ii), (iii) and (iv), as well as strictly positive reals $r_{1},\ldots,r_{m}$ such that $r_{1}\boldsymbol{s}_{1}+\cdots+r_{m}\boldsymbol{s}_{m}=\boldsymbol{0}$. Indeed, this is true if and only if the largest linear subspace $\mathcal{L}$ of the $\mathbb{R}_{\geq 0}$-cone $\langle\mathcal{S}\rangle_{\mathbb{R}_{\geq 0}}$ contains some cell $\operatorname{Cell}(\boldsymbol{d}_{i*})$ for each $i\in J_{+}$, some cell $\operatorname{Cell}(\boldsymbol{d}_{*j})$ for each $j\in J_{-}$, and some cell $\operatorname{Cell}(\boldsymbol{e}_{k})$ for each $k\in J_{0}$. This is decidable in PTIME since the number of cells in $\mathcal{L}$ is finite (at most nine).
@@
First, we decide in PTIME which of the six cases is true for $H$, using Lemma 14. In cases (ii), (iii) and (v), the Group Problem for $\langle\mathcal{G}\rangle$ has positive answer. In cases (i), (iv) and (vi), Proposition 15, Corollary 19 and Corollary 25 show the required PTIME decidability result.
@@ 6. Extensions and obstacles to Semigroup Membership
Under the additional assumption that $G$ is a group, we can decide Semigroup Membership for $\langle\mathcal{G}\rangle$ in all except one cases. Recall Theorem 5, if $G$ contains a non-abelian free subgroup, then $\langle\mathcal{G}\rangle$ is a group by Proposition 6. Hence Semigroup Membership reduces to Group Membership, and is decidable by the result of Delgado [DR17]. If $G$ is virtually solvable, then consider the six cases in Proposition 7. Case (ii), (iii) and (v) are easy since $\langle\mathcal{G}\rangle$ becomes a group. In case (i), deciding whether $(I,\boldsymbol{a})\in\langle\mathcal{G}\rangle$ reduces to solving the linear equation $n_{1}\boldsymbol{a}_{1}+n_{2}\boldsymbol{a}_{2}+\cdots n_{K}\boldsymbol{a}_{K}=\boldsymbol{a}$ for $(n_{1},\ldots,n_{K})\in\mathbb{Z}_{\geq 0}^{K}\setminus\{\boldsymbol{0}\}$, and is decidable by integer programming. In case (iv), Semigroup Membership for $\langle\mathcal{G}\rangle$ reduces to Semigroup Membership in the Heisenberg group $\operatorname{H}_{3}(\mathbb{Q})$, which is decidable by the result of Colcombet et al. [COSW19]. Only case (vi) remains unsolved.

### v6

@@ 5.6. $H$ is generated by a positive scale $A$
    Hence, in both cases, $\boldsymbol{c}_{1}+C_{1}\boldsymbol{c}_{M}\in\operatorname{Cell}(\boldsymbol{d}_{ij})\subseteq\mathcal{S}$. We then show that $\boldsymbol{c}_{1}+C_{1}\boldsymbol{c}_{M}+C_{1}\boldsymbol{x}^{\prime}$ is an $\mathbb{R}_{>0}$-combination of elements in $\mathcal{S}$. Since $\boldsymbol{x}^{\prime}$ is either zero or an $\mathbb{R}_{>0}$-combination of elements in $\mathcal{S}$, write $\boldsymbol{x}^{\prime}=\sum_{i=1}^{m}r_{i}\boldsymbol{s}_{i}$, where $m\geq 0,r_{i}>0,\boldsymbol{s}_{i}\in\mathcal{S}$. Then $C_{1}\boldsymbol{x}^{\prime}=\sum_{i=1}^{m}r_{i}C_{1}\boldsymbol{s}_{i}$ is still an $\mathbb{R}_{>0}$-combination of elements in $\mathcal{S}$ by Equation (8). Hence, $\boldsymbol{c}_{1}+C_{1}\boldsymbol{c}_{M}+C_{1}\boldsymbol{x}^{\prime}$ is an $\mathbb{R}_{>0}$-combination of elements in $\mathcal{S}$.
  3. (3)
    $t_{1}t_{M}=0$. Suppose $t_{1}=0$, the case where $t_{M}=0$ can be done analogously.
@@
    where $\boldsymbol{y}$ and $\boldsymbol{y}^{\prime}$ can be written as an $\mathbb{R}_{>0}$-linear combination of elements in $\mathcal{S}$. Therefore $\boldsymbol{x}=\boldsymbol{y}+\boldsymbol{y}^{\prime}$ also satisfies this claim.
@@
The semigroup $\langle\mathcal{G}\rangle$ is a group if and only if there exists a finite set of vectors $\{\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{m}\}\subseteq\mathcal{S}$ that satisfies the four conditions (i)-(iv) in Lemma 23 with $\boldsymbol{x}=\boldsymbol{0}$.
@@
Therefore, $\boldsymbol{s}_{t}$ can always be written as a strictly positive linear combination of vectors $\boldsymbol{v}$ with $(I,\boldsymbol{v})\in\langle\mathcal{G}^{\prime}\rangle$. Hence, by condition (i), there exist strictly positive reals $r_{1},\ldots,r_{M}$ and vectors $\boldsymbol{v}^{\prime}_{1},\ldots,\boldsymbol{v}^{\prime}_{M}$, such that $r_{1}\boldsymbol{v}^{\prime}_{1}+\cdots+r_{M}\boldsymbol{v}^{\prime}_{M}=\boldsymbol{0}$ and $(I,\boldsymbol{v}^{\prime}_{t})\in\langle\mathcal{G}^{\prime}\rangle$ for $t=1,\ldots,M$. Furthermore, conditions (ii), (iii) and (iv) show that every letter $(A^{\prime}_{i},\boldsymbol{a}^{\prime}_{i})$ in $\mathcal{G}^{\prime}$ appears at least once in a word representing $(I,\boldsymbol{v}^{\prime}_{t})$ for some $t\in\{1,\ldots,M\}$.
@@
Next, suppose $\langle\mathcal{G}\rangle$ is a group. We show that there exists an $\mathbb{R}_{>0}$-linear combination of elements in $\mathcal{S}$ equal to $\boldsymbol{0}$, that satisfies the conditions (ii), (iii) and (iv).
@@
Then by Lemma 23, the vector $\boldsymbol{0}$ can be written as an $\mathbb{R}_{>0}$-linear combination of a finite number of elements in $\mathcal{S}$, satisfying the conditions (ii), (iii) and (iv).
@@
If the generator $A$ of the group $\langle A_{1},\ldots,A_{K}\rangle\cong\mathbb{Z}$ is a positive scale, it is decidable in PTIME whether $\langle\mathcal{G}\rangle$ is a group.
@@
Since $\langle A_{1},\ldots,A_{K}\rangle\cong\mathbb{Z}$, we have $J_{+}\neq\emptyset,J_{-}\neq\emptyset$. Given $\mathcal{S}$ as a set of cells, it is decidable whether there exist a finite non-empty set of vectors $\{\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{m}\}\subseteq\mathcal{S}$ satisfying conditions (ii), (iii) and (iv), as well as strictly positive reals $r_{1},\ldots,r_{m}$ such that $r_{1}\boldsymbol{s}_{1}+\cdots+r_{m}\boldsymbol{s}_{m}=\boldsymbol{0}$. Indeed, this is true if and only if there exists a linear subspace $\mathcal{L}$ contained in the $\mathbb{R}_{\geq 0}$-cone generated by $\mathcal{S}$, with the property that $\mathcal{L}$ intersects some cell $\operatorname{Cell}(\boldsymbol{d}_{i*})$ for each $i\in J_{+}$, some cell $\operatorname{Cell}(\boldsymbol{d}_{*j})$ for each $j\in J_{-}$, and some cell $\operatorname{Cell}(\boldsymbol{e}_{k})$ for each $k\in J_{0}$. This is decidable in PTIME since the number of cells in $\mathcal{L}$ is finite (at most nine).
@@
First, we decide in PTIME which of the six cases is true for $H$, using Lemma 14. In cases (ii), (iii) and (v), the Group Problem for $\langle\mathcal{G}\rangle$ has a positive answer. In cases (i), (iv) and (vi), Proposition 15, Corollary 19 and Corollary 25 show the required PTIME decidability result.
@@ 6. Extensions and obstacles to Semigroup Membership
Under the additional assumption that $G$ is a group, we can decide Semigroup Membership for $\langle\mathcal{G}\rangle$ in all cases except one. Recall Theorem 5. If $G$ contains a non-abelian free subgroup, then $\langle\mathcal{G}\rangle$ is a group by Proposition 6. Hence Semigroup Membership reduces to Group Membership, and is decidable by the result of Delgado [DR17]. If $G$ is virtually solvable, then consider the six cases in Proposition 7. Case (ii), (iii) and (v) are easy since $\langle\mathcal{G}\rangle$ becomes a group. In case (i), deciding whether $(I,\boldsymbol{a})\in\langle\mathcal{G}\rangle$ reduces to solving the linear equation $n_{1}\boldsymbol{a}_{1}+n_{2}\boldsymbol{a}_{2}+\cdots n_{K}\boldsymbol{a}_{K}=\boldsymbol{a}$ for $(n_{1},\ldots,n_{K})\in\mathbb{Z}_{\geq 0}^{K}\setminus\{\boldsymbol{0}\}$, and is decidable by integer programming. In case (iv), Semigroup Membership for $\langle\mathcal{G}\rangle$ reduces to Semigroup Membership in the Heisenberg group $\operatorname{H}_{3}(\mathbb{Q})$, which is decidable by the result of Colcombet et al. [COSW19]. Only case (vi) remains unsolved.